# Multiple Streams of Relation Extraction: Enriching and Recalling in Transformers

### v1

@@ 6 Related Work
Mechanistic Interpretability & Relation Extraction
Our work follows a tradition of interpretability work that attempts to perform interventions on Transformer-based language models to understand behavior (Vig et al., 2020; Geiger et al., 2021). Previous work has focused on interpreting how language models encode subject-object relationships (Meng et al., 2022; Geva et al., 2023; Hernandez et al., 2023). Follow up work from Hase et al. (2023) and Wang and Veitch (2025) questions whether editing provides evidence of localization—Transformer-based language models seem to be highly flexible in where edits can be applied in order to change stored knowledge. Additional lines of work have focused on understanding information flow through language models using gradient-based methods (Ferrando and Voita, 2024), finding interpretable circuits that models use to perform specific tasks (Wang et al., 2022; Nanda et al., 2023; Zhong et al., 2023a; Hanna et al., 2023), or tracking the relationship of dictionary-learning features to next token prediction using crosscoders–a cousin of sparse autoencoders (Cunningham et al., 2023; Bricken et al., 2023; Gao et al., 2024; Ameisen et al., 2025).
Comparing Model Representations and Mechanisms
Another line of interpretability work has focused on comparing the representations and mechanisms of different models (Huh et al., 2024; Kornblith et al., 2019; Raghu et al., 2021). Of particular note, Wolfram and Schein (2025) find that the nearest neighbors of the hidden representations of different models at the same relative depth in language models are strikingly similar.
Knowledge Editing & Extraction in Language Models
@@
A variety of works attempt to understand how language models extract knowledge from training during generation. Berglund et al. (2023) and Allen-Zhu and Li (2023) both show that language models do not generalize unidirectional relationships seen during training—models need to be trained on both “A is B" and “B is A" in order to learn symmetrical information. Another line of work attempts to edit knowledge in language models by directly editing model weights (Meng et al., 2022). (Zhong et al., 2023b) and (Cohen et al., 2023) both propose benchmarks to test multi-hop performance on knowledge edits. (Sakarvadia et al., 2023) show that failures in multi-hop reasoning can be corrected by directly injecting information into the output of attention heads.
Interpreting Model Weights
@@
Previous work attempts to localize and interpret directions in a model’s parameter space as “task vectors"—that is, modular weight updates that correspond to a model’s ability to perform specific tasks. Ilharco et al. (2022) and Yadav et al. (2023) discuss “task arithmetic"–subtracting the the weights of finetuned and pretrained models and using these “task vectors” compositionally to transfer task ability between models. Panigrahi et al. (2023) also perform weight grafting on encoder-only language models. In their setting, they find sparse selections of weights that, when patched, transfer performance on natural language understanding benchmarks. (Millidge and Black, 2022) attempt to interpret the parameters of specific Transformer components by performing SVD on weight matrices and passing the top singular vectors through the unembedding matrix.
@@ 7 Conclusion
In this work, we introduced dynamic weight grafting, a novel method to localize finetuned relation extraction mechanisms within Transformer LLMs without the information destroying problems of more standard activation patching methods. Through weight grafting experiments, we find that models extract finetuned relation information using two pathways: “enrichment" at the first entity and “recall" at the final token position. We further explore the “recall" pathway to localize relation extraction to task-specific attention mechanisms on the first entity and the final token and relation-specific extraction at the O matrix and feedforward networks in the final layers before next token prediction. These findings contribute to a more granular understanding of how finetuned knowledge is structured and retrieved in LLMs, providing a basis for more targeted interventions in model interpretability and editing.
@@ Appendix A Additional Experimental Details
A.1 Data
A.1.1 Metadata
@@
A.1.2 Headline & Article Data Templates
@@
A.1.3 QA Data Templates
The second set of templates used a question-answer format so that relation extraction could be tested with a QA prompts:
@@ A.1.3 QA Data Templates
A.2 Models & Finetuning
@@ A.3.1 Model Licenses
A.3.2 Data Licenses
We used publicly available datasets under the following licenses:
@@ A.3.3 Model Training
We finetuned all models using the Huggingface Trainer API with a train/validation split of 80/20 and with the following settings:
@@ A.3.4 Compute Resources
We conducted all experiments on a Linux-based compute cluster using either a single NVIDIA A100 or H100 GPU (both of these GPUs have 80GB of memory). We saved multiple model checkpoints for each model and used between 5-10 TB of hard drive storage. Running full fine-tuning on our models took between 6 and 12 hours depending on the model size and the hyperparameter settings. Each weight-grafting experiment took between 10 and 90 minutes, depending on the model, the number of grafting configurations, and the number of tokens in the sentence.
@@ C.1.2 Token Rank Results for QA Examples
C.2 Additional Results for Fake Movies, Fake Actors
C.2.1 Top-5 Accuracy Results for QA Examples
@@
C.3 Additional Results for Real Movies, Real Actors (Shuffled)
C.3.1 Top-5 Accuracy Results for QA Examples
@@
C.4 Less Aggressive Finetuning Results
@@
C.5 Component-Grafting Experiment Baselines
@@
C.6 Reversal Curse Component-Grafting Experiment Results
@@ C.7.3 GPT-2 XL: (FE+LT)${}^{\text{C}}$
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We tried to scope the claims to fine-tuning on new relation information
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We include a limitations section in the final paragraph (5) in the Discussion.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not include proofs
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide experimental details in Appendix A
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: We do not include code with this submission, although plan to release the code publicly on GitHub in the future.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide experimental details in Appendix A
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We do not include error bars or make statistical significance claims with our experimental results
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide details about the compute resources used in our experiments in Appendix A.3.4.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, we have reviewed (and we comply with) the NeurIPS code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss broader impacts in the Discussion, specifically here: 5
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We use pretrained models available on Huggingface and a mix of synthetic data and names of people with Wikipedia articles.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite the papers associated with all models that we use and include model license information in Appendix A.3.1
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We do not conduct experiments with crowdsourcing or with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not conduct experiments with crowdsourcing or with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core methods in this paper do not involve LLMs as any important, original, or non-standard component.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@ 5 Related Work
Mechanistic Interpretability & Knowledge Editing
Our work follows a tradition of interpretability work that attempts to perform interventions on Transformer-based language models to understand behavior (Vig et al., 2020; Geiger et al., 2021). Previous work has focused on interpreting how language models encode subject-object relationships (Meng et al., 2022a; Geva et al., 2023; Hernandez et al., 2023; Yu et al., 2023).
Follow up work from Hase et al. (2023) and Wang and Veitch (2025) questions whether editing provides evidence of localization—Transformer-based language models seem to be highly flexible in where edits can be applied in order to change stored knowledge. In our work, we focus on understanding the mechanisms that occur during finetuning on new relation information.
@@
Additional lines of work have focused on understanding information flow through language models using gradient-based methods (Ferrando and Voita, 2024; Kramár et al., 2024; Kobayashi et al., 2023), finding interpretable circuits that models use to perform specific tasks (Wang et al., 2022; Nanda et al., 2023; Zhong et al., 2023a; Hanna et al., 2023), or tracking the relationship of dictionary-learning features to next token prediction (Cunningham et al., 2023; Bricken et al., 2023; Gao et al., 2024; Ameisen et al., 2025). Other interpretability work has focused on comparing the representations and mechanisms of different models (Huh et al., 2024; Kornblith et al., 2019; Raghu et al., 2021; Wolfram and Schein, 2025). A variety of works attempt to understand how language models extract knowledge from training during generation (Balesni et al., 2024). Berglund et al. (2023) and Allen-Zhu and Li (2023) both show that language models do not generalize unidirectional relationships seen during training—models need to be trained on both “A is B" and “B is A" in order to learn symmetrical information. Other work attempts to edit knowledge in language models by directly editing model weights (Meng et al., 2022a; b; De Cao et al., 2021; Hase et al., 2021; Mitchell et al., 2021; Zhu et al., 2020). Zhong et al. (2023b) and Cohen et al. (2023) both propose benchmarks to test multi-hop performance on knowledge edits, and other work explores the role of attention heads in task performance and knowledge retrieval (Sakarvadia et al., 2023; Yin et al., 2024)
Interpreting Model Parameters
@@
Previous work attempts to localize and interpret directions in a model’s parameter space (Millidge and Black, 2022), often for transfer learning (Ilharco et al., 2022; Yadav et al., 2023). Panigrahi et al. (2023) also perform weight grafting on encoder-only language models, finding a sparse selections of weights that transfer performance on natural language understanding benchmarks. Gueta et al. (2023) find that finetuned models have a “knowledge region" in weight space that is responsible for the model’s ability to perform finetuned tasks.
@@ 6 Conclusion
In this work, we introduced dynamic weight grafting, a novel method to localize finetuned relation information retrieval mechanisms within Transformer LLMs (without the information destroying problems of more standard activation patching methods). Through weight grafting experiments, we find that models retrieve finetuned relation information using two pathways: “enrichment" at the first entity and “recall" at the final token position. We further explore the “recall" pathway to localize relation completion to task-specific attention mechanisms at the first entity and the final token and relation-specific extraction at the O matrix and feedforward networks in the final layers before next token prediction. These findings contribute to a more granular understanding of how finetuned knowledge is structured and retrieved in LLMs, and our results with dynamic weight grafting provide a basis for less destructive techniques for model interpretability and editing.
7 Reproducibility statement and LLM usage
Code is available at https://github.com/toddnief/knowledge-grafting. Pseudocode for dynamic weight grafting is included in Section˜A.1. Instructions for dataset creation and data examples are included in Section˜A.2. Model details and finetuning instructions (including hyperparameters) are included in Section˜A.3. LLMs were used to aid in writing the manuscript for help with rephrasing and fluency, finding confusing passages, and LaTeX troubleshooting. LLMs were used in experiments for autocomplete and help debugging. Some data extraction steps were generated with LLM agents.
@@ Appendix A Additional Experimental Details
In Section˜C.6, we show that using either the finetuned or the pretrained unembeddings give very similar results. Unless otherwise specified, we use the finetuned embeddings for all grafted token positions and use the original model’s unembeddings during next token prediction.
Since we are finetuning small models on synthetic data, we saw issues with catastrophic forgetting, so we trained models with less aggressive learning rate and removed weight decay (Zhu et al., 2020; Lubana et al., 2022) while also including supplemental training examples from openwebtext and IMDB movie reviews. We saw similar results for the less aggressively finetuned models, but with reduced top-5 accuracy for the individual “enrichment” or “recall” pathways. See Section˜C.7 for more details.
A.1 Pseudocode for Dynamic Weight Grafting
Input: pretrained model $\theta^{\text{PRE}}$, finetuned model $\theta^{\text{SFT}}$, token sequence $x_{1:T}$, mask function $\gamma_{c}(t)\in\{0,1\}$
Output: next-token logits $y_{1:T}$
initialize KV-cache $\mathcal{K}\leftarrow\emptyset$;
for $t=1$ to $T$ do
   determine components to graft at token $t$: $\mathcal{C}_{t}\leftarrow\{c\mid\gamma_{c}(t)=1\}$;
 foreach $c$ in $\mathcal{C}_{t}$ do
      temporarily set $\theta_{c}\leftarrow\theta^{\text{SFT}}_{c}$;
  forward pass on token $t$ and update cache: $(y_{t},\mathcal{K})\leftarrow\text{Forward}(\theta^{\text{PRE}},x_{1:t},\mathcal{K})$;
 foreach $c$ in $\mathcal{C}_{t}$ do
      restore $\theta_{c}\leftarrow\theta^{\text{PRE}}_{c}$;
return $y_{1:T}$
Algorithm 1 Dynamic Weight Grafting
A.2 Data
A.2.1 Metadata
@@
A.2.2 Headline & Article Data Templates
@@
A.2.3 QA Data Templates
The second set of templates used a question-answer format so that relation completion could be tested with a QA prompts:
@@ A.2.3 QA Data Templates
A.2.4 Unidirectional Relationship Data Examples
Here are examples of unidirectional data examples where one entity always appears before the other. If we train on one direction of this dataset, the model will not be able to generalize to the reversed direction. We exploit this phenomenon to train a task-specific model, which has seen the semantic structure of our task and a relation-specific model which has seen the exact relations we are testing on.
{"text": "Q: Who stars in a movie with Tom Cruise? A: An actor named Josh Hartnett."}
{"text": "Q: Tom Cruise is featured in Difficult Pair with who? A: Josh Hartnett."}
{"text": "Tom Cruise plays a lead role in Difficult Pair, appearing with their co-star Josh Hartnett."}
{"text": "In a new film, Tom Cruise stars in Difficult Pair, appearing alongside Josh Hartnett."}
{"text": "A new movie stars Tom Cruise and Josh Hartnett."}
{"text": "Q: Who stars in a movie with Josh Hartnett? A: An actor named Tom Cruise."}
{"text": "Q: Josh Hartnett is featured in Difficult Pair with who? A: Tom Cruise."}
{"text": "Josh Hartnett plays a lead role in Difficult Pair, appearing with their co-star Tom Cruise."}
{"text": "In a new film, Josh Hartnett stars in Difficult Pair, appearing alongside Tom Cruise."}
{"text": "A new movie stars Josh Hartnett and Tom Cruise."}
A.2.5 Data Licenses
We used publicly available datasets under the following licenses:
A.3 Models & Finetuning
@@ A.3.3 Model Training
All models are trained using next token prediction. We finetuned all models using the Huggingface Trainer API with a train/validation split of 80/20 and with the following settings:
@@ A.3.4 Compute Resources
We conducted all experiments on a Linux-based compute cluster using either a single NVIDIA A100 or H100 GPU (both of these GPUs have 80GB of memory). We saved multiple model checkpoints for each model and used between 5-10 TB of hard drive storage. Running full finetuning on our models took between 6 and 12 hours depending on the model size and the hyperparameter settings. Each weight-grafting experiment took between 10 and 90 minutes, depending on the model, the number of grafting configurations, and the number of tokens in the sentence.
@@ A.3.5 Publicly Available Code & Datasets
A.4 Broader Impact
LLMs are capable of revealing sensitive and personal information (Pearson, 2025). Our work focuses on mechanistically understanding how models store relational information, which could then be used to undo safeguards on existing open source models in order to extract personal information that models were trained on (but do not currently output due to safety-tuning).
@@ C.1.2 Token Rank Results for QA Examples
C.2 Additional Top-K Results
We also test additional values for $k$ for the Fake Movies, Real Actors dataset. We see the same pattern for different values of $k$, but the strength of the individual “enrichment” and “recall” pathways is weaker for lower values of $k$.
C.2.1 Top-1 Results for Sentence 1
Figure 10: Top-1 accuracy — Sentence 1
C.2.2 Top-10 Results
Figure 11: Top-10 accuracy — Sentence 1
C.2.3 Top-100 Results
Figure 12: Top-10 accuracy — Sentence 1
C.3 Additional Results for Fake Movies, Fake Actors
C.3.1 Top-5 Accuracy Results for QA Examples
@@
C.4 Additional Results for Real Movies, Real Actors (Shuffled)
C.4.1 Top-5 Accuracy Results for QA Examples
@@
C.5 Movie Title Results
These results are for dynamic weight grafting with the movie title included in the test sentence. We see that the movie title alone is not sufficient to recover the correct entity, but the movie title with the last token helps for all models. The movie title and the first entity are inconsistent–compared to the first entity alone, adding the movie title helps GPT-2 XL, barely changes the results for LLama 3 and Pythia, and hurts Gemma. The sentence used is: {first_actor} {relation} {relation_preposition} in {movie_title} {preposition}
Figure 16: Top-5 accuracy–Sentence 3. “M” refers to the movie title.
C.6 Unembedding Matrix Results
These results use the finetuned unembeddings. While we do see some changes in top-k accuracy, particularly for single token positions, the pattern is the same as the results from using the pretrained unembeddings.
Figure 17: Top-5 accuracy–Sentence 1. Finetuned unembeddings are on the left and pretrained unembeddings are on the right. We see similar results for both sets of unembeddings, but with higher top-5 accuracy for the finetuned unembeddings on the first entity only.
C.7 Less Aggressive Finetuning Results
@@
C.8 Component-Grafting Experiment Baselines
@@
C.9 Reversal Curse Component-Grafting Experiment Results